# GARCH copulas and GARCH-mimicking copulas

### v1

@@ 5.2 Simulation study of jointly symmetric case
Copula $p$ | $(\epsilon_{t})$ Gauss t4 t2.5 Gauss t4 t2.5
$C^{t}_{\nu}$ (iv-ast) 1 482 0 115 56 0 77
mix-t 2 484 2 117 54 2 79
mix-Clayton 1 1630 426 178 444 284 127
mix-Gumbel 1 867 150 171 181 113 138
mix-Joe 1 1936 593 265 619 424 214
iv-Clayton-180 1 483 146 281 127 98 198
iv-Gumbel 1 1121 444 368 375 300 272
iv-Joe 1 604 50 134 97 58 113
mix-K-Clayton 3 1124 430 91 338 281 52
mix-K-Gumbel 3 728 154 55 152 117 46
mix-K-Joe 3 1634 597 167 585 428 133
iv-K-Clayton-180 3 0 147 145 42 91 101
iv-K-Gumbel 3 552 448 210 283 296 141
iv-K-Joe 3 47 54 0 0 56 0
Table 2: Comparison of copula fits when jointly symmetric copulas are fitted to simulated data from symmetric ARCH(1) and GARCH(1,1) processes by the PML method. Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC.
5.3 Simulation study of h-symmetric case with leverage
@@
Copula $p$ | $(\epsilon_{t})$ Gauss t4 t2.5 Gauss t4 t2.5
iv-ast 1 980 224 222 683 262 173
iv-Clayton-180 1 1009 358 363 806 384 299
iv-Joe 1 1114 267 237 727 319 200
iva-ast 2 258 0 131 23 0 68
iva-Clayton-180 2 364 152 280 209 139 202
iva-Joe 2 374 48 152 54 60 100
iv-K-Clayton-180 3 655 361 239 715 386 238
iv-K-Joe 3 672 271 113 619 323 132
iva-K-Clayton-180 4 22 153 131 151 136 112
iva-K-Joe 4 0 52 0 0 61 0
Table 3: Comparison of AIC values when various h-symmetric copulas are fitted to simulated data from ARCH(1) and GARCH(1,1) processes with symmetric innovations and leverage by the PML method. Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC.
In Table 3 we include a selection of the copulas from Table 2, omitting the mixture copulas, since the inverse-v-transformed copulas typically do better, and also omitting the iv-Gumbel copula, since iv-Clayton-180 and iv-Joe generally do better. We add copulas with the prefix ‘iva’ (referring to asymmetric inverse v-transform) which are models of the form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),|2v-1|)$ where $\mathcal{V}_{1}$ is a member of the family $\mathcal{V}(u;\delta,\kappa)$ in (13) with $\delta=0.5$, so that only one extra unknown parameter $\kappa$ is introduced.
The first panel contains jointly symmetric and exchangeable copulas; the second panel contains h-symmetric copulas based on an exchangeable copula for $c^{*}$ but incorporating an asymmetric v-transform for the first variable; the third panel contains jointly symmetric non-exchangeable copulas using the Khoudraji construction; the final panel contains h-symmetric copulas based on a non-exchangeable $c^{*}$ and incorporating an asymmetric v-transform for the first variable.
For the two models with t4 innovations the best fitting copula is the absolute spherical t copula incorporating an asymmetric v-transform. For the other models the best fitting copula is a Khoudraji-transformed version of the Joe copula incorporating an asymmetric v-transform. Results are very similar for the ARCH(1) and GARCH(1,1) process.
5.4 Simulation study of case of asymmetric innovations
Model ARCH(1) GARCH(1,1)
Copula $p$ leverage leverage
iv-ast 1 168 679 49 436
iv-Clayton-180 1 376 881 208 673
iv-Joe 1 220 728 92 487
ivl-ivl-ast 3 116 201 23 92
ivl-ivl-Clayton-180 3 347 558 189 403
ivl-ivl-Joe 3 164 231 64 140
ivl-ivl-K-Clayton-180 5 336 561 191 403
ivl-ivl-K-Joe 5 168 235 68 144
iv2-iv2-ast 5 0 0 0 0
iv2-iv2-Clayton-180 5 228 293 159 277
iv2-iv2-Joe 5 47 24 45 50
iv2-iv2-K-Clayton-180 7 221 297 160 279
iv2-iv2-K-Joe 7 50 28 49 54
Table 4: Comparison of AIC values when various copulas are fitted to simulated data from ARCH(1) and GARCH(1,1) processes with asymmetric innovations by the PML method. Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC.
@@ 5.4 Simulation study of case of asymmetric innovations
The first set of 3 models in Table 4 consists of jointly symmetric and exchangeable copulas, included as a benchmark. The iv-ast copula is the best of these but substantial improvements can be made by introducing asymmetry.
@@
The second set of 3 models is based on copula densities of the form $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ where $\mathcal{V}(\cdot;\delta)$ is the linear v-transform with fulcrum at $\delta$ (asymmetric when $\delta\neq 0.5$) and $c^{*}$ is either the ast, Clayton survival or Joe copula [^4] It should be noted that full maximum likelihood estimation with respect to all parameters including the parameters of the v-transform presents no problems for these models since the ast, Clayton survival and Joe copula densities do not have asymptotes at $(0,0)$. If this were not the case we can run into a threshold estimation issue, as discussed in Bladt and McNeil, 2022b .. These obviously lower the AIC values considerably at the expense of 2 extra parameters. The third set of 2 models applies the Khoudraji transformation to $c^{*}$ in the case of Clayton and Joe, but this only offers an improvement in one case (Clayton-180 and ARCH(1) process).
@@
The fourth set of models is based on copula densities $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ where $\mathcal{V}(\cdot;\delta,\kappa)$ is the 2-parameter v-transform in (13); such v-transforms are both non-linear and asymmetric. These models give the best results of all, with the model where $c^{*}$ is the ast copula taking first place for all 4 datasets, followed by the model based on Joe. Addition of the Khoudraji transformation in the final set of models again offers no discernible improvement to the models based on Joe. It seems that allowing asymmetric v-transform largely removes the need to consider the Khoudraji transformation, even when the v-transforms are linear.
6 Mimicking GARCH(1,1) processes with simplified d-vines
6.1 Practical issues in the construction of simplified d-vines
In the final sections of this paper we provide some guidance to the construction of simplified d-vine copula models to mimic the behaviour of GARCH-type processes and offer potential alternative models for financial time series. The previous sections have given some insight into the kinds of pair copulas we should consider. We now turn to some of the issues that arise when we want to embed these in simplified vine structures.
Equation (1) in the Introduction implies that the density $c_{(n)}$ of the copula of the variables $(X_{1},\ldots,X_{n})^{\top}$ in a strictly stationary time series model, such as GARCH, can be written in the form
$c_{(n)}(u_{1},\ldots,u_{n})=\\ \prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}(B_{k-1}(u_{j-k};\bm{u}_{[j-k+1:j-1]}),R_{k-1}(u_{j};\bm{u}_{[j-1:j-k+1]})\mid F_{X}^{-1}(u_{j-k+1}),\ldots,F_{X}^{-1}(u_{j-1}))$ (22)
where, for $k\in\mathbb{N}$, we define the functions $R_{k}:(0,1)\times(0,1)^{k}\to(0,1)$ and $B_{k}:(0,1)\times(0,1)^{k}$
$R_{k}(x;\bm{u})$ $=$ $F_{X_{k+1}|\bm{X}_{[1:k]}}(F_{X}^{-1}(x)\mid F_{X}^{-1}(u_{k}),\ldots,F_{X}^{-1}(u_{1}))$
$B_{k}(x;\bm{u})$ $=$ $F_{X_{1}|\bm{X}_{[2:k+1]}}(F_{X}^{-1}(x)\mid F_{X}^{-1}(u_{1}),\ldots,F_{X}^{-1}(u_{k}))$
which we refer to as forward and backward Rosenblatt functions, respectively. The notation is based on the one used in Bladt and McNeil, 2022a with minor changes. Note that, in (22), we adopt the convention that $R_{0}(x;\cdot)=B_{0}(x;\cdot)=x$ for all $x\in(0,1)$.
In practice we want to mimic the copula (22) with a simplified d-vine, i.e. a model where the copula has joint density of the simpler form
$c_{(n)}(\bm{u})=\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}\Big{(}B_{k-1}(u_{j-k};\bm{u}_{[j-k+1:j-1]}),R_{k-1}(u_{j};\bm{u}_{[j-1:j-k+1]})\Big{)}.$ (23)
In this case the Rosenblatt functions are determined in a recursive, interlacing fashion by $R_{1}(x;u)=h_{1,1}(u,x)$, $B_{1}(x;v)=h_{1,2}(x,v)$ and
[…]
where $h_{k,1}(u,v)=\frac{\partial}{\partial u}C_{k}(u,v)$ and $h_{k,2}(u,v)=\frac{\partial}{\partial v}C_{k}(u,v)$ are the so-called $h$-functions of the copula $C_{k}$ with density $c_{k}$.
It seems unlikely that we can perfectly mimic a GARCH-type process with a simplified d-vine, since the true copula (22) cannot be shown to conform to the simplifying assumption. However, the weak dependence of $c_{2}(u,v;w)$ on the conditioning variable investigated in Section 3.5 gives some optimism that we could at least approximate a GARCH-type model with a simplified d-vine.
Other than ARCH(1)-type processes, which are Markov processes of order 1, GARCH(1,1)-type processes have infinite memory in the sense that $X_{t}$ depends on all previous values $X_{t-k}$, for $k\in\mathbb{N}$, albeit the serial dependence on past observations dies away relatively quickly; for example, the autocorrelation function of the squared values of the classical GARCH(1,1) process decays like an ARMA(1,1) process. To mimic a GARCH(1,1)-type process, it seems natural to consider d-vine processes with infinite-order partial copula dependence as in Bladt and McNeil, 2022a . These are models which are built from a sequence of pair copulas $(C_{k})_{k\in\mathbb{N}}$ which is parsimoniously parameterized using the Kendall partial autocorrelation function (kpacf) or, in other words, the sequence of Kendall rank correlations of $(C_{k})_{k\in\mathbb{N}}$.
GARCH(1,1) models are very parsimoniously parameterized. For example, a model that incorporates both leverage and skewed Student t innovations would have only 6 parameters ($\alpha_{0}$, $\alpha_{1}$, $\beta_{1}$, $\gamma_{1}$, $\nu$, $\lambda$). In our d-vine analogues we also need to control the number of parameters and the method of parameterization via the kpacf developed in Bladt and McNeil, 2022a can be adapted to our context by applying it to an inverse-v-transformed version of the copula series $(C^{*}_{k})_{k\in\mathbb{N}}$. To be more precise, each pair copula $C_{k}$ used in the sequence defining the d-vine would be constructed to have density $c_{k}(u,v)=c_{k}^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ where the v-transforms $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ would be two different, probably asymmetric functions with a fixed form for all $k$; in essence they would take care of the effects caused by leverage and asymmetric innovations in a similar manner to $\gamma_{1}$ and $\lambda$ in the aforementioned GARCH(1,1) model. The copula sequence $(C^{*}_{k})_{k\in\mathbb{N}}$ would contain Joe, Clayton survival or ast copulas with Kendall correlations $\tau_{k}$ that coincided with the terms of the kpacf of a classical Gaussian ARMA(1,1) process, or some other low-order linear process. The combination of ARMA(1,1) and two linear v-transforms would give precisely 4 parameters to describe the d-vine copula process, although two or more parameters would also be required to model the marginal distribution of $(X_{t})_{t\in\mathbb{Z}}$.
We have listed the ast copula $C^{|t|}_{\nu}$ as a candidate pair copula for the sequence $(C^{*}_{k})_{k\in\mathbb{N}}$, but it is not clear whether there is a simple parametric formula to link Kendall’s tau to the degree of freedom parameter $\nu$. Simple calculations from (20) show that
$\tau(\nu)=64\int_{0}^{0.5}\int_{0}^{0.5}C_{\nu}^{t}(u,v)c_{\nu}^{t}(u,v)\mathrm{d}u,\mathrm{d}v-1$ (25)
where $C_{\nu}^{t}$ and $c_{\nu}^{t}$ are the spherical t copula and its density, and that $\tau(\nu)$ ranges in the interval $(0,1)$ as $\nu$ ranges in the interval $(0,\infty)$. An analytical expression for $\tau(\nu)$ would make the ast copula a very attractive choice for these models but, in the absence of such an expression, the Joe copula may be the easier option.
Finally we consider the choice of v-transforms for $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$. To estimate a d-vine model by maximum likelihood we have to be able to evaluate a likelihood based on (23) which means evaluating the Rosenblatt forward and backward functions using the recursive expressions (24). This in turn entails evaluating multiple $h$-functions. Moreover, to generate realizations from a copula with joint density (23), we have to be able to invert the Rosenblatt forward functions which entails the evaluation of inverse $h$-functions. To this end it is convenient to have closed form expressions for the copulas $C_{k}$ and their $h$-functions and inverse $h$-functions. For copula densities $c_{k}$ obtained from the inverse-v-transformation construction in (15) this is unfortunately not straightforward, except in the case where we have copula densities of the form $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ constructed from two linear v-transforms $\mathcal{V}_{\delta_{1}}$ and $\mathcal{V}_{\delta_{2}}$ and a copula density $c^{*}$ which itself has a tractable joint distribution function and $h$-functions. We investigate this in the next section.
@@ 6.1 Practical issues in the construction of simplified d-vines
6.2 Tractable processes using linear v-transforms
@@
$C(u,v)=\delta_{1}^{I_{\{u\leqslant\delta_{1}\}}}(\delta_{1}-1)^{I_{\{u>\delta_{1}\}}}\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}C^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))+\delta_{1}v+\delta_{2}u-\delta_{1}\delta_{2}$ (26)
@@
$\mathcal{V}\left(h_{1}(u,v);\delta_{2}\right)$ $=h_{1}^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ (27)
@@
First note that $\frac{\partial^{2}}{\partial u\partial v}C(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$. We just need to check that the expression (26) defines a copula, which entails checking that $C(0,v)=0$ and $C(1,v)=v$ for any $v$; by symmetry it would follow that $C(u,0)=0$ and $C(u,1)=u$ for any $u$. Now
$C(0,v)=\delta_{1}\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}\mathcal{V}(v;\delta_{2})+\delta_{1}v-\delta_{1}\delta_{2}$
and this yields 0 for both $v\leqslant\delta_{2}$ and $v>\delta_{2}$; moreover
$C(1,v)=(\delta_{1}-1)\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}\mathcal{V}(v;\delta_{2})+\delta_{1}v+\delta_{2}(1-\delta_{1})$
and this yields $v$ for both $v\leqslant\delta_{2}$ and $v>\delta_{2}$. The $h$-functions are obtained by differentiation of (26) and the identities (27) are easy to verify. Noting that $h_{1}(u,\delta_{2})=\delta_{2}$ and $h_{2}(u,\delta_{1})=\delta_{1}$, it is also straightforward to derive the inverses of the $h$-functions. ∎
We end by considering the special case of the model with linear v-transforms in which the two v-transforms are identical. We find that the resulting d-vine copula also belongs to the class of vt-d-vines in Bladt and McNeil, 2022b ; these models have been used with real financial data and found to be tractable and effective. The statement of the result and proof use the notation $\mathcal{V}(\bm{u};\delta)$ for the element-by-element v-transformation of a vector $\bm{u}$ with the same linear v-transform.
@@
Suppose we construct a d-vine copula (23) using a sequence of copulas $(C_{k})_{k\in\mathbb{N}}$ whose densities are all of the form $c_{k}(u,v)=c_{k}^{*}(\mathcal{V}(u;\delta),\mathcal{V}(v;\delta))$ for a sequence of copulas $(C_{k}^{*})_{k\in\mathbb{N}}$ with densities $(c_{k}^{*})_{k\in\mathbb{N}}$ and a single linear v-transform $\mathcal{V}(\cdot,\delta)$. Then
$c_{(n)}(\bm{u})=\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}^{*}\Big{(}B^{*}_{k-1}(\mathcal{V}(u_{j-k};\delta);\mathcal{V}(\bm{u}_{[j-k+1:j-1]};\delta)),R^{*}_{k-1}(\mathcal{V}(u_{j};\delta);\mathcal{V}(\bm{u}_{[j-1:j-k+1]};\delta))\Big{)}$ (28)
where $(R_{k}^{*})_{k\in\mathbb{N}}$ and $(B_{k}^{*})_{k\in\mathbb{N}}$ are the forward and backward Rosenblatt functions corresponding to the copula sequence $(C_{k}^{*})_{k\in\mathbb{N}}$.
@@
$\mathcal{V}\left(B_{k-1}(u;\bm{v});\delta\right)=B^{*}_{k-1}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v};\delta))\quad\text{and}\quad\mathcal{V}\left(R_{k-1}(u;\bm{v});\delta\right)=R^{*}_{k-1}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v};\delta))$ (29)
hold for any $k\in\mathbb{N}$, $u\in(0,1)$ and $\bm{v}\in(0,1)^{k-1}$. We will argue by induction. The equalities (29) are obviously true for $k=1$ since, by our convention, $B_{0}(u;\cdot)=B_{0}^{*}(u;\cdot)=u$ and $R_{0}(u;\cdot)=R_{0}^{*}(u;\cdot)=u$ for all $u$. Assume the equalities (29) are true when $k=l\in\mathbb{N}$ and consider $k=l+1$. For the first equality we can show, by first using formula (27) in Proposition 7 and then applying the inductive assumption, that

### v2

@@ 5 Conclusion
Copula (form of density) $p$ | $(\epsilon_{t})$ Gauss t4 t2.5 Gauss t4 t2.5
$c^{|t|}_{\nu}(|2u-1|,|2v-1|)$ 1 1068 224 223 720 264 187
$c_{\theta}^{\text{J}}(|2u-1|,|2v-1|)$ 1 1202 267 238 764 321 214
$c_{\theta}^{\text{Cl},180}(|2u-1|,|2v-1|)$ 1 1097 358 364 843 386 313
$c^{|t|}_{\nu}(\mathcal{V}(u;0.5,\kappa),|2v-1|)$ 2 346 0 132 60 2 82
$c_{\theta}^{\text{J}}(\mathcal{V}(u;0.5,\kappa),|2v-1|)$ 2 462 48 153 91 62 114
$c_{\theta}^{\text{Cl},180}(\mathcal{V}(u;0.5,\kappa),|2v-1|)$ 2 452 152 281 246 141 216
$c_{\nu,a_{1},a_{2}}^{|t|}(|2u-1|,|2v-1|)$ 3 678 227 113 614 267 131
$c_{\theta,a_{1},a_{2}}^{\text{J}}(|2u-1|,|2v-1|)$ 3 760 271 114 656 325 146
$c_{\theta,a_{1},a_{2}}^{\text{Cl},180}(|2u-1|,|2v-1|)$ 3 743 361 240 752 388 252
$c_{\nu,a_{1},a_{2}}^{|t|}(\mathcal{V}(u;0.5,\kappa),|2v-1|)$ 4 0 2 0 0 0 0
$c_{\theta,a_{1},a_{2}}^{\text{J}}(\mathcal{V}(u;0.5,\kappa),|2v-1|)$ 4 88 52 1 37 63 14
$c_{\theta,a_{1},a_{2}}^{\text{Cl},180}(\mathcal{V}(u;0.5,\kappa),|2v-1|)$ 4 110 153 132 188 138 126
Table S1: Comparison of AIC values when various h-symmetric copulas are fitted to simulated data from ARCH(1) and GARCH(1,1) processes with symmetric innovations and leverage by the PML method. Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC.
In Table S1 we include a selection of the copulas from Table 2, omitting the mixture copulas, since the inverse-v-transformed copulas typically do better, and also omitting the inverse-v-transformed Gumbel copula, since the other inverse-v-transformed copulas do better. We add copulas of the form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),|2v-1|)$ where $\mathcal{V}_{1}$ is a member of the family $\mathcal{V}(u;\delta,\kappa)$ in (12) with $\delta=0.5$, so that only one extra unknown parameter $\kappa$ is introduced.
The first panel contains jointly symmetric and exchangeable copulas; the second panel contains h-symmetric copulas based on an exchangeable copula for $c^{*}$ but incorporating an asymmetric v-transform $\mathcal{V}(u;0.5,\kappa)$ for the first variable; the third panel contains jointly symmetric non-exchangeable copulas using the Khoudraji construction (17); the final panel contains h-symmetric copulas based on a non-exchangeable $c^{*}$ and incorporating an asymmetric v-transform for the first variable.
The best results in all cases are obtained from copulas developed from the absolute spherical t copula $C^{|t|}_{\nu}$ that are not jointly symmetric and that incorporate the non-symmetric v-transform, as theory suggests. In 5 of 6 cases the inverse v-transformation of the Khoudraji non-exchangeable extension of $C^{|t|}_{\nu}$ is selected while in one case (ARCH(1) with t4 innovations) the inverse v-transformation of $C^{|t|}_{\nu}$ is selected. Models based on the Joe copula also do well.
S5.2 Simulation study of case of asymmetric innovations
@@
Copula (form of density) $p$ | leverage absent present absent present
$c^{|t|}_{\nu}(|2u-1|,|2v-1|)$ 1 169 679 49 436
$c_{\theta}^{\text{J}}(|2u-1|,|2v-1|)$ 1 221 728 92 487
$c_{\theta}^{\text{Cl},180}(|2u-1|,|2v-1|)$ 1 377 881 208 673
$c^{|t|}_{\nu}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ 3 117 201 23 92
$c^{\text{J}}_{\nu}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ 3 165 231 64 140
$c^{\text{Cl},180}_{\nu}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ 3 348 558 189 403
$c^{|t|}_{\nu,a_{1},a_{2}}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ 5 116 205 25 94
$c^{\text{J}}_{\nu,a_{1},a_{2}}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ 5 169 235 68 144
$c^{\text{Cl},180}_{\nu,a_{1},a_{2}}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ 5 337 561 191 403
$c^{|t|}_{\nu}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ 5 1 0 0 0
$c^{\text{J}}_{\nu}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ 5 48 24 45 50
$c^{\text{Cl},180}_{\nu}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ 5 229 293 159 277
$c^{|t|}_{\nu,a_{1},a_{2}}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ 7 0 4 1 4
$c^{\text{J}}_{\nu,a_{1},a_{2}}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ 7 51 28 49 54
$c^{\text{Cl},180}_{\nu,a_{1},a_{2}}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ 7 222 297 160 279
Table S2: Comparison of AIC values when various copulas are fitted to simulated data from ARCH(1) and GARCH(1,1) processes with asymmetric innovations by the PML method. Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC.
@@
The first set of 3 models in Table S2 consists of jointly symmetric and exchangeable copulas, included as a benchmark. The spherical t copula is the best of these but substantial improvements can be made by introducing asymmetry.
@@
The second set of 3 models is based on copula densities of the form $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ where $\mathcal{V}(\cdot;\delta)$ is the linear v-transform with fulcrum at $\delta$ (asymmetric when $\delta\neq 0.5$) and $c^{*}$ is either the ast, Clayton survival or Joe copula. It should be noted that full maximum likelihood estimation with respect to all parameters including the parameters of the v-transform presents no problems for these models since the ast, Clayton survival and Joe copula densities do not have asymptotes at $(0,0)$. If this were not the case we can run into a threshold estimation issue, as discussed in Bladt and McNeil, 2022b . Adding the parameters $\delta_{1}$ and $\delta_{2}$ lowers the AIC values considerably. The third set of 3 models applies the Khoudraji transformation to $c^{*}$ before application of the inverse linear v-transform, but this barely offers any improvement on the previous set of models; the AIC is only really lowered for the Clayton survival copula in the ARCH(1) model without leverage, and survival Clayton is the worst of the three candidates in this case.
@@
The fourth set of models is based on copula densities $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ where $\mathcal{V}(\cdot;\delta,\kappa)$ is the 2-parameter v-transform in (12); such v-transforms are both non-linear and asymmetric. These models give the best results of all, with the model where $c^{*}$ is the ast copula best for 3 out of 4 datasets datasets, followed by the model based on Joe. Addition of the Khoudraji transformation in the final set of models only offers an improvement for the ARCH(1) model without leverage, and this improvement is very minor.
We conclude that allowing asymmetric v-transform largely removes the need to consider the Khoudraji transformation, even when the v-transforms are linear. This has positive consequences for practical work: evaluating the density of Khoudraji-extended copulas of type (17) is computationally challenging since the density involves the copula $C_{\theta}$ and its h-functions, as well as is density $c_{\theta}$. This is particularly burdensome when $C_{\theta}=C^{|t|}_{\nu}$, the ast copula, as numerical integration is required to evaluate $C^{|t|}_{\nu}$.
@@
Appendix S6 D-vine processes using pair copulas based on linear v-transforms
S6.1 Calculating $h$-functions
@@
$C(u,v)=\delta_{1}^{I_{\{u\leqslant\delta_{1}\}}}(\delta_{1}-1)^{I_{\{u>\delta_{1}\}}}\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}C^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))+\delta_{1}v+\delta_{2}u-\delta_{1}\delta_{2}$ (S4)
@@
$\mathcal{V}\left(h_{1}(u,v);\delta_{2}\right)$ $=h_{1}^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ (S5)
@@
The joint density $c$ of the copula $C$ defined in  (S4) can be easily shown to satisfy $c(u,v)=\frac{\partial^{2}}{\partial u\partial v}C(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$. We just need to check that it is a valid copula, which entails checking that $C(0,v)=0$ and $C(1,v)=v$ for any $v$, which is straightforward. By symmetry, it then follows that $C(u,0)=0$ and $C(u,1)=u$ for any $u$. The $h$-functions are obtained by differentiation of (S4) and the identities (S5) are easy to verify. Noting that $h_{1}(u,\delta_{2})=\delta_{2}$ and $h_{2}(u,\delta_{1})=\delta_{1}$, it is also straightforward to derive the inverses of the $h$-functions. ∎
S6.2 Link to vt-D-vine models
In the special case of a model with two identical linear v-transforms we now show that the resulting D-vine copula also belongs to the class of vt-D-vines in Bladt and McNeil, 2022b which has been applied successfully to real financial data. The use of two different linear v-transforms does however give much more flexibility.
For this result we introduce sequences of conditional distribution functions (Rosenblatt functions) $B_{k}:(0,1)\times(0,1)^{k}\to(0,1)$ and $R_{k}:(0,1)\times(0,1)^{k}\to(0,1)$ defined by
$\begin{aligned} B_{k-1}(x;\bm{u})&=\mathbb{P}(U_{1}\leqslant x\mid U_{2}=u_{1},\ldots,U_{k}=u_{k-1})=F_{U_{1}\mid\bm{U}_{[2:k]}}(x\mid\bm{u}_{[1:k-1]}),\\ R_{k-1}(x;\bm{u})&=\mathbb{P}(U_{k}\leqslant x\mid U_{k-1}=u_{1},\ldots,U_{1}=u_{k-1})=F_{U_{k}\mid\bm{U}_{[1:k-1]}}(x\mid\bm{u}_{[k-1:1]}),\end{aligned}\quad k\in\mathbb{N}.$ (S6)
Note the reversal of the order of the conditioning values $\bm{u}_{[k-1:1]}=(u_{k-1},\ldots,u_{1})^{\top}$ in the second equation. Note also that we interpret $B_{0}(x;\cdot)=R_{0}(x;\cdot)=x$. The Rosenblatt functions satisfy $R_{1}(x;u)=h_{1,1}(u,x)$, $B_{1}(x;v)=h_{1,2}(x,v)$ and
$\begin{aligned} R_{k}(x;\bm{u})&=h_{k,1}\left(B_{k-1}(u_{k};\bm{u}_{[k-1:1]}),R_{k-1}(x;\bm{u}_{[1:k-1]})\right),\\ B_{k}(x;\bm{u})&=h_{k,2}\left(B_{k-1}(x;\bm{u}_{[1,k-1]}),R_{k-1}(u_{k};\bm{u}_{[k-1:1]})\right),\end{aligned}\quad k\geqslant 2,$ (S7)
where $h_{k,1}(u,v)=\frac{\partial}{\partial u}C_{k}(u,v)$ and $h_{k,2}(u,v)=\frac{\partial}{\partial v}C_{k}(u,v)$ are the $h$-functions of the copula $C_{k}$ with density $c_{k}$. This allows us to write the density in (18) as
$c_{U_{1},\ldots,U_{n}}(u_{1},\ldots,u_{n})=\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}\Big(B_{k-1}(u_{j-k};\bm{u}_{[j-k+1:j-1]}),R_{k-1}(u_{j};\bm{u}_{[j-1:j-k+1]})\Big).$ (S8)
@@
Suppose we construct a d-vine copula (S8) using a sequence of copulas $(C_{k})_{k\in\mathbb{N}}$ whose densities are all of the form $c_{k}(u,v)=c_{k}^{*}(\mathcal{V}(u;\delta),\mathcal{V}(v;\delta))$ for a sequence of copulas $(C_{k}^{*})_{k\in\mathbb{N}}$ with densities $(c_{k}^{*})_{k\in\mathbb{N}}$ and a single linear v-transform $\mathcal{V}(\cdot,\delta)$. Then the joint density $c_{U_{1},\ldots,U_{n}}(u_{1},\ldots,u_{n})$ may be written as
$\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}^{*}\Big(B^{*}_{k-1}(\mathcal{V}(u_{j-k};\delta);\mathcal{V}(\bm{u}_{[j-k+1:j-1]};\delta)),R^{*}_{k-1}(\mathcal{V}(u_{j};\delta);\mathcal{V}(\bm{u}_{[j-1:j-k+1]};\delta))\Big)$ (S9)
where $(R_{k}^{*})_{k\in\mathbb{N}}$ and $(B_{k}^{*})_{k\in\mathbb{N}}$ are the forward and backward Rosenblatt functions for the copula sequence $(C_{k}^{*})_{k\in\mathbb{N}}$ and where we use the compact notation $\mathcal{V}(\bm{u};\delta)=(\mathcal{V}(u_{1};\delta),\ldots,\mathcal{V}(u_{d};\delta))^{\top}$.
@@
$\mathcal{V}\left(B_{k-1}(u;\bm{v});\delta\right)=B^{*}_{k-1}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v};\delta))\quad\text{and}\quad\mathcal{V}\left(R_{k-1}(u;\bm{v});\delta\right)=R^{*}_{k-1}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v};\delta))$ (S10)
hold for any $k\in\mathbb{N}$, $u\in(0,1)$ and $\bm{v}\in(0,1)^{k-1}$. We will argue by induction. The equalities (S10) are obviously true for $k=1$ since, by our convention, $B_{0}(u;\cdot)=B_{0}^{*}(u;\cdot)=u$ and $R_{0}(u;\cdot)=R_{0}^{*}(u;\cdot)=u$ for all $u$. Assume the equalities (S10) are true when $k=l\in\mathbb{N}$ and consider $k=l+1$. For the first equality we can show, by first using formula (S5) in Proposition S3 and then applying the inductive assumption, that